# Product Design Using a Generative Adversarial Network: Incorporating Consumer Preferences and External Data

### v2

@@
Abstract
The development of generative artificial intelligence (AI) enables large-scale product design automation. However, this automated process usually does not incorporate consumer preference information from the internal dataset of a company. Furthermore, external sources such as social media and user-generated content (UGC) websites often contain rich product design and consumer preference information, but such information is not utilized by companies when generating designs. We propose a semi-supervised deep generative framework that integrates consumer preferences and external data into the product design process, allowing companies to generate consumer-preferred designs in a cost-effective and scalable way. We train a predictor model to learn consumer preferences and use predicted popularity levels as additional input labels to guide the training procedure of a continuous conditional generative adversarial network (CcGAN). The CcGAN can be instructed to generate new designs with a certain popularity level, enabling companies to efficiently create consumer-preferred designs and save resources by avoiding the development and testing of unpopular designs. The framework also incorporates existing product designs and consumer preference information from external sources, which is particularly helpful for small or start-up companies that have limited internal data and face the “cold-start” problem. We apply the proposed framework to a real business setting by helping a large self-aided photography chain in China design new photo templates. We show that our proposed model performs well in terms of generating appealing template designs for the company.
@@ Abstract
1 Introduction
Product design is pivotal for firms to attract consumers and achieve success. The traditional product design process involves hiring professional human designers to produce new designs in small batches, which is a costly and unscalable approach. The rapid development of generative artificial intelligence (AI) can potentially make the task of generating new product aesthetic designs effortless. For instance, the big generative adversarial network (BigGAN) (Brock et al.,, 2018) is capable of generating high-resolution and diverse images across multiple classes. Platforms such as Artbreeder enable users to create unique artworks by blending existing images. [^1] See https://www.artbreeder.com/. From product packaging to car components and retail displays, industrial designers can use generative AI to explore more ideas and develop initial design concepts significantly faster than they could with traditional methods (McKinsey & Company,,2023).
While these AI tools are used either before the design generation stage to analyze consumer data and extract consumer needs or after the design generation stage to predict consumer demand for new designs, most of the current AI applications do not systemically incorporate large-scale consumer preference information into the design generation process. During the design generation stage, consumer preferences are integrated in a limited and ad hoc manner. For example, generative AI-based text-to-image software can be used to design new clothing of a certain style through iterative prompting (McKinsey & Company,, 2023). However, these prompts are not generated in a systematic way and may not capture consumers’ heterogeneous tastes. Even if new designs are generated automatically and at a low cost, how attractive they are to consumers remains unknown. To predict consumer demand for new designs, firms often need to employ A/B tests or “theme clinics” by asking consumers to evaluate different aesthetic designs. Theme clinics are costly; for example, firms usually spend more than \$100,000 conducting a single vehicle design test when designing cars (Burnap et al.,, 2023). In general, the task of predicting consumer demand is performed after generating new designs, which potentially wastes the resources spent developing and testing unpopular designs.
@@ 1 Introduction
In contrast to the lack of incorporating consumer preferences into generative AI applications, rich consumer preference information exists in user-generated content (UGC) and remains underutilized by companies in practice. By nature, UGC data inherently contain rich consumer preference information as users voluntarily generate the associated content. For example, social media and UGC websites contain vast sets of photos taken by individual users in front of various backgrounds. The fact that a user chose to take a photo in front of a particular background means that the user liked the background. As such, UGC images can be used to understand how consumers evaluate new photo background designs. The backgrounds in UGC images can also serve as useful inputs for training generative models to generate new photo template designs. Despite the wide availability of UGC data, their usage in marketing is mostly descriptive and focuses on extracting consumer preference information or predicting consumer demand. Prescriptive UGC usage (i.e., designing products for companies) is scant both in marketing academia and in practice. No clear guidance has been provided on how to use unstructured UGC data to directly make unstructured firm decisions such as product design.
Another challenge lies in the cost of finding a scalable solution for incorporating consumer preferences and leveraging UGC data. The development of a generative AI algorithm may be costly and unaffordable for many companies. Training an algorithm also requires a large training dataset, which may not be available for small companies with limited internal data or start-up companies facing a “cold-start” problem.
@@
Our study fills these gaps and addresses three key questions. (1) How can consumer preference information be incorporated into the product design process, and how can generative AI be guided to generate new designs that are appealing to consumers? (2) How can firms leverage the rich design and preference information contained in UGC during the design generation process, especially for firms facing “cold-start” problems? (3) How can the generation process be automated in a systematic, scalable, and cost-efficient way?
We propose a semi-supervised deep generative framework that creates consumer-preferred product designs in an automatic and cost-effective manner. The framework consists of a predictor model that learns consumer preferences and a generative model that uses the outputs of the prediction model to guide the generation process. This study has several highlights. First, different from previous studies in which consumer preference information was only used for training a predictor model, we incorporate consumer preference information into both our predictor and generator models. In particular, the predicted popularity for each product serves as a label and an additional input to the generator, which allows the generator to create products that are more likely to be preferred by consumers. Second, we leverage UGC acquired from external data sources. The various product designs contained in this UGC enrich the generative model and allow it to incorporate new product features that are not included in the internal data of firms. The rich consumer preference information embedded in the UGC improves the supervision of the generative model and allows it to generate appealing designs that go beyond the scope of the internal preferences and designs.
We present different versions of the generative model to illustrate the advantages of our proposed framework. We start with a simple GAN that does not incorporate any consumer preferences. We then gradually add consumer preferences to the generative process, first adding only internal preferences and then adding both internal and external preferences.
Specifically, the baseline model uses a deep convolutional GAN (DCGAN). Internal designs and external designs are used as inputs to generate new designs without incorporating consumer preferences. This approach extends the current industry practice by incorporating new designs from UGC. The second model uses the continuous conditional GAN (CcGAN) introduced by Ding et al., (2020). Compared to traditional GANs, the CcGAN can directionally generate new images of certain types. In particular, the CcGAN takes image labels as additional generative model inputs, which enables it to generate new images with a specific label. For example, a traditional GAN can use a set of dog images to produce a new set of dog images, while the CcGAN can instruct its model to create images of a specific type of dog (e.g., a golden retriever). The CcGAN is well suited for our task because it allows us to incorporate consumer preferences and generate new product designs that are more likely to be chosen by consumers. We first train a predictor model to learn consumer preferences from consumers’ historical choice data regarding existing designs. We then use the predicted choice probabilities acquired from the predictor model as additional input labels to guide the training process of the CcGAN. Finally, we instruct the trained CcGAN to generate new product designs that have higher predicted choice probabilities. We examine two versions of the CcGAN: an enhanced model and an advanced model. The enhanced model only uses internal consumer preference information to train the predictor model. The advanced model uses both internal and external consumer preference information to train the predictor model, and this variant represents the final version of our proposed framework.
We apply our proposed framework to a real business setting: a self-aided photography chain in China. The company operates self-aided photo booths in large shopping malls in major cities. The photo booths allow consumers to choose photo templates, take photos with the chosen templates, and print their photos. These photo templates are the main product offerings of the company, so it is crucial for the firm to design appealing photo templates to attract consumers. However, as a start-up company, the firm has limited resources and does not have a systematic way to design new templates. The firm also faces a “cold-start” problem with limited internal data. Our proposed framework can help the firm generate new consumer-preferred photo template designs in an automatic and cost-effective manner.
@@
The internal dataset contains 2195 consumers who had taken at least one photo between October 2017 and August 2018. For each consumer, we obtain her full order history, including her selected photo templates and the photos taken with these chosen templates. We further collect external images from UGC websites. These external images are of similar types to those of the internal images, with humans taking photos in front of various backgrounds. The external dataset serves two purposes. First, it provides training inputs for the constructed generative models, allowing them to incorporate new product features derived from external product designs. It addresses the issue that the internal data contain limited style variations. Second, the rich history of external preferences inherent in the external data allows the predictor model to recognize popular product features from the external sources and guide the generative model to produce “out-of-the-box” appealing designs that do not exist in the internal data.
Comparing the performances of the generated templates produced by different models, we find that the advanced model with both internal and external consumer preferences performs best. The new designs generated by the advanced model are predicted to be more likely to be preferred by consumers by both the predictor model and outside metrics. Our results suggest that merely incorporating internal consumer preferences is not sufficient; it is important to further incorporate external consumer preferences, as external preferences provide additional information on how consumers evaluate external designs, thereby guiding the generative model to incorporate the popular features of the external designs that do not exist in the internal designs. Overall, the results highlight the importance of incorporating consumer preferences into the generation process.
@@
Managerially, our work provides an automated product design solution for firms, which is especially valuable for smaller start-up firms that face cold-start problems and budget constraints. Firms have internal data on their existing product designs and the choices of consumers among the existing products. Our framework enables firms to further incorporate product design ideas and consumer tastes from external websites in a systematic, scalable, and cost-effective way. The proposed design process can also dynamically evolve as more internal and external data are incorporated as inputs, which enables the firms to rapidly respond to consumer demand shifts.
@@
Our proposed model enhances the traditional product design process in three ways. First, it offers a superior alternative to the prevailing industry practice that relies solely on pure GAN models. The traditional methods do not consider consumer tastes during the design generation process; thus, generating a large set of new designs and screening unpopular designs on an ex post facto basis are needed. Our method generates new designs that are preferred by consumers ex ante, enabling companies to spend less time and resources on new designs that are later shown to be less appealing. Second, our proposed model provides an effective solution for product design ideation. The traditional product design ideation process relies on professionally trained human designers, who have limited capacity and may not systematically learn the dynamically evolving requirements of a large set of consumers. Our method leverages design ideas from UGC that reflect the real-time preferences of a large user base. Third, our proposed model offers a more cost-effective solution for obtaining consumer evaluations of new product designs. In contrast to traditional methods such as conducting costly consumer surveys, we leverage UGC as a source of real-time, self-revealed consumer preferences that are relevant and dynamically updating. Our method is applicable to other product design settings (e.g., video meeting backgrounds) where consumer preference information widely exists and external resources (e.g., UGC) are easily accessible.
2 Conceptual Model for Generating New Designs
2.1 Overview of the Conceptual Model
We propose a semi-supervised deep generative framework for generating consumer-preferred new product designs in a systematic and cost-efficient way. Our proposed model has three unique advantages. First, it allows firms to leverage product designs from external sources such as UGC on social media and image websites. This is especially beneficial to start-up firms with limited internal product designs. Second, it incorporates consumer preference information from the internal data of a firm during the design generation process so that the generated new designs are preferred by consumers. Third, it incorporates consumer preference from external sources, which allows firms to leverage the rich and widely available consumer preference information contained in UGC.
To illustrate these three advantages, we introduce three versions of the generative model. The first is a baseline model that uses product designs from both internal and external sources to generate new designs. This approach illustrates the first advantage of leveraging external product designs from UGC. The second variant builds on the baseline model and further incorporates consumer preference information acquired from the internal data of a firm. Compared with the first model, this version illustrates the second advantage of allowing internal consumer preferences to guide the generative process and directionally generate more appealing designs. The third variant is the full version that incorporates product design and consumer preference information from both internal and external sources. This method demonstrates all three advantages of our proposed framework.
In this subsection, we describe the intuitions and workflows of the three models. We discuss the implementation details of the models in the following subsections.
@@ 2.1 Overview of the Conceptual Model
Studies have explored the use of UGC in marketing. Most of these studies were descriptive and focused on analyzing UGC to extract information and gain insights for firm operations. Some of them have analyzed textual content. For example, Netzer et al., (2012) extracted brand associations from UGC on forums. Culotta and Cutler, (2016) extracted the social connections of brands on Twitter. Other works analyzed visual content (Zhang and Luo,, 2023; Zhang et al.,, 2017; Pavlov and Mizik,, 2019) . For instance, Liu et al., (2020) analyzed consumer-generated images on social media to study the brand perceptions of consumers. We contribute to the literature by going one step further, from descriptive to prescriptive work. Our generative model not only extracts information from unstructured UGC but also uses this information to make unstructured product design decisions for firms.
Following industry standards, we use GANs (Goodfellow et al.,, 2014) to generate new product designs in the baseline model. We choose GAN models over other generative models such as fully visible belief networks (FVBNs) (Frey,, 1998) and variational autoencoders (VAEs) (Kingma and Welling,, 2013; Rezende et al.,, 2014). The reason for this is that GANs were designed to avoid the many disadvantages associated with other generative models and are regarded as producers of higher-quality samples than those of other methods. As we focus on image synthesis-based tasks, we use the DCGAN (Radford et al.,, 2015). The DCGAN is considered one of the most important steps forward in terms of designing and training stable GANs and has become the de facto standard given its high-level performance in image tasks.
@@
We first train a predictor model to learn the internal preferences of consumers, which we incorporate into the generative process. The predictor model utilizes the historical photos taken by internal consumers as inputs. It learns a mapping between a consumer’s appearance (i.e., her facial traits) and her preferred product designs, assuming that consumers with similar appearances will prefer similar product designs. To address the high dimensionality of the target problem, we preprocess the observed consumer faces and product designs using embedding models, obtaining low-dimensional vector representations for both consumer appearances and product designs. Using these embeddings as inputs, we train a deep prediction model to capture high-level abstractions in the prediction task. We call this predictor model with internal preferences “Prediction Model 1”, as illustrated in Figure 1. The trained predictor model can be used to predict the popularity of any given product design. Specifically, for each pair consisting of a product design and an internal consumer, the trained predictor calculates the probability that this consumer will choose the corresponding product. By averaging over the selection probabilities among all internal consumers, we can obtain a popularity measure of the particular product.
@@
The trained predictor model is further integrated into the generator model in a semi-supervised deep learning framework. We employ a conditional GAN model (Mirza and Osindero,, 2014), which extends the traditional GAN by enabling the conditional generation of outputs. Intuitively, the traditional GAN can use a set of dog images to produce a new set of dog images, but one cannot instruct the model to create images of a specific type of dog. The conditional GAN allows one to condition the network on additional information such as class labels. During training, the conditional GAN takes images with labels (bulldogs, golden retrievers, etc.) as inputs and learns the differences between the images with different labels. In this way, it gains the ability to generate new images with a specific label (a specific type of dog). The conditional GAN is well suited for our task because it allows us to incorporate consumer preferences and generate new product designs that are more likely to be chosen by consumers. In our context, we use the popularity measures derived from the predictor model as additional inputs (i.e., class labels of product designs). The predicted popularity levels guide the training procedure of the conditional GAN so that internal consumer preferences are incorporated into the generative model. Once the generative model is trained, we can instruct the model to generate new product designs with higher popularity levels. In this way, the conditional GAN can generate new designs that are more appealing to consumers.

### v3

@@
Product Design Using a Generative Adversarial Network: Incorporating Consumer Preferences and External Data
Hui Li ,  Jian Ni , Fangzhu Yang HKU Business School, University of Hong Kong, Hong Kong. Email: huil1@hku.hk. Pamplin College of Business, Virginia Tech. Email: jiann@vt.edu. Economics Division, Bates White. Email: Fangzhu517@hotmail.com.
The rise of generative artificial intelligence (AI) has facilitated automated product design but often neglects valuable consumer preference data within companies’ internal datasets. Additionally, external sources such as social media and user-generated content (UGC) platforms contain substantial untapped information on product design and consumer preferences, yet remain underutilized. We propose a novel framework that transforms the product design paradigm to be data-driven, automated, and consumer-centric. Our method employs a semi-supervised deep generative architecture that systematically integrates multidimensional consumer preferences and heterogeneous external data. The framework is both generative and preference-aware, enabling companies to produce consumer-aligned designs with enhanced cost efficiency. Our framework trains a specialized predictor model to comprehend consumer preferences and utilizes predicted popularity metrics to guide a continuous conditional generative adversarial network (CcGAN). The trained CcGAN can directionally generate consumer-preferred designs, circumventing the expenditure associated with testing suboptimal candidates. Using external data, our framework offers particular advantages for start-ups or other resource-constrained companies confronting the “cold-start” problem. We demonstrate the framework’s efficacy through an empirical application with a self-operated photography chain, where our model successfully generated superior photo template designs. We also conduct web-based experiments to verify our method and confirm its effectiveness across varying design contexts.
@@
Introduction
Product design is crucial for firms to attract consumers and succeed. Traditionally, this involves hiring professional designers to create small batches of new designs, which is costly and lacks scalability. To evaluate attractiveness of new designs, companies need to gather consumer feedback through focus groups, surveys, and market tests, such as A/B tests or theme clinics. Theme clinics are usually expensive, with vehicle design tests costing over \$100,000 (Burnap et al., 2023). After collecting feedback, firms analyze it using techniques such as conjoint analysis (Green and Srinivasan, 1990), choice modeling (McFadden, 1986), and preference mapping (Urban and Hauser, 1993). The insights derived from the analyses are then used to refine designs to maximize market share or profits (Allenby and Rossi, 1998). This iterative, manual process results in significant delays from concept development to market launch.
The rapid advancement of generative AI has the potential to significantly accelerate the product design process. In the design generation phase, generative AI enables industrial designers to quickly explore a wider array of ideas and develop initial concepts much faster than traditional methods (McKinsey & Company,, 2023). For instance, generative AI text-to-image tools like DALL-E, Stable Diffusion, and Midjourney can be used to design clothing of a specific style through iterative prompting. [^1] https://openai.com/index/dall-e-3/,https://stability.ai/news/stable-diffusion-3,https://www.midjourney.com/home
However, a significant gap exists in current AI applications for product design: the design generation stage lacks systematic integration of large-scale consumer preferences (Burnap et al., 2023). AI tools are typically used either before design generation to analyze consumer data and identify needs or after to predict demand for new designs (Dzyabura and Hauser, 2011; Toubia et al., 2003). During design generation, consumer preferences are either missing or incorporated in a limited, ad hoc manner. Although text-to-image tools enable design generation through iterative prompting, these prompts are not systematically generated and may fail to capture diverse consumer tastes. Additionally, the attractiveness of the generated designs to consumers remains uncertain. Predicting consumer demand occurs after design creation, potentially leading to wasted resources on developing and testing unpopular designs.
@@ Introduction
In contrast to the lack of incorporating consumer preferences into generative AI applications, rich consumer preference information exists in user-generated content (UGC) and remains underutilized by companies in practice. UGC naturally contains valuable consumer insights, as users voluntarily create content. For instance, social media and UGC sites host numerous user-taken photos in various settings, indicating user preferences for those backgrounds. These images can help understand consumer evaluations of new photo backgrounds and serve as inputs for training generative models to create new photo background designs. Despite the wide availability of UGC data, their usage in marketing is mostly descriptive and focuses on extracting preferences or predicting demand. Prescriptive usage of UGC for product design is scant both in marketing academia and in practice. There is no clear guidance on using unstructured UGC data for direct unstructured product designs.
Additionally, finding a scalable solution to integrate consumer preferences and leverage UGC data poses challenges. Developing a generative AI algorithm can be expensive, potentially unaffordable for many companies. Training such algorithms requires large datasets, which may be unavailable to small or startup companies facing a “cold-start” problem. Furthermore, even after training, generating new designs can be costly due to the high time costs and prices of generative AI tools (see the literature review for a detailed discussion).
@@
Our study addresses existing gaps by focusing on three key questions: (1) How can consumer preference information be integrated into the product design process to guide generative AI in creating appealing designs? (2) How can firms utilize the rich design and preference data in UGC during the design generation stage, particularly for those facing “cold-start” issues? (3) How can the design generation process be automated systematically, scalably, and cost-effectively?
We propose a novel framework to transform the product design process in the era of generative AI. Our approach utilizes a semi-supervised deep generative framework based on Generative Adversarial Network (GAN) to automatically and cost-effectively create consumer-preferred product designs. Compared to other generative models such as diffusion models (see the literature review for a detailed comparison), our model offers several advantages: 1) it generates realistic images; 2) it is prompt-free, eliminating the need for textual prompts; 3) it incorporates large-scale consumer preferences; 4) it leverages UGC; 5) it enables fast and cost-effective design generation; [^2] Our trained model can generate a new design in just 0.01 seconds using a consumer-grade laptop, whereas state-of-the-art diffusion models require at least 10 seconds and rely on 2 to 3 enterprise-grade GPUs. 6) it is tailored for personalized design generation.
To illustrate the practical application of our framework, we conduct an empirical study with a self-operated photography chain in China. This company operates photo booths in major city shopping malls, allowing customers to select templates, take photos, and print them. As the photo templates are key products, it is crucial for the company to design appealing ones to attract customers. Traditionally, this would involve hiring professional designers and conducting costly consumer tests for each design. As a start-up with limited resources and internal data, the company faces a “cold-start” problem. Our framework enables the company to systematically integrate revealed preferences from both internal consumer choices and external UGC sources, facilitating rapid and concurrent design generation and evaluation. This approach transforms the process from a sequential design-then-test method to an integrated design-with-preferences strategy, fundamentally changing how firms can develop consumer-oriented products in the era of generative AI.
@@
The self-operated photography chain provides us internal dataset that contains 2195 consumers who had taken at least one photo between October 2017 and August 2018. For each consumer, we observe her full order history, including her selected photo templates and the photos taken with these chosen templates. The mapping from internal consumers to their chosen templates represents internal consumer preference. We further collect external images from UGC websites. These external images are of similar types to those of the internal images, with humans taking photos in front of various backgrounds. They represent both existing external designs and external consumer preferences.
We apply our proposed framework to this empirical context. The framework includes a predictor model that learns consumer preferences and a generative model that uses these predictions to guide design generation. We present variations of the generative model to highlight the framework’s benefits, starting with a deep convolutional GAN (DCGAN) that generates new designs from internal and external existing designs in UGC without incorporating consumer preferences. Next, we integrate consumer preferences into the generative process by adapting continuous conditional GAN (CcGAN) from Ding et al., (2020) to our context. Unlike traditional GANs, CcGAN can directionally generate specific types of images by using image labels as additional inputs to train the generator, making it ideal for incorporating consumer preferences. We first train a predictor model with historical consumer choice data to learn preferences, then use these predicted choice probabilities as additional input labels to train the CcGAN. Importantly, we modify the original CcGAN in Ding et al., (2020) to accommodate the ordinal nature of consumer preference labels. The trained CcGAN is then instructed to generate new product designs with higher choice probabilities. We explore two CcGAN versions: an enhanced model using only internal consumer preferences for training the predictor and an advanced model utilizing both internal and external preferences, representing our final framework version. We show that the advanced model can be easily adapted to create personalized designs for individual consumers.
The performance of the generated templates is evaluated through three methods. First, a distance-based metric is employed to assess the similarity of the generated templates to popular existing ones. Popular templates share common features that consumers favor. We use the Caffe BVLC reference model (Jia et al.,,2014) to obtain embeddings for both existing and generated templates. We calculate the distance between generated templates and the most popular and unpopular existing templates. A smaller (larger) distance to the most popular (unpopular) templates suggests higher potential popularity. Second, consumer evaluations are gathered through an online survey where participants choose between existing and newly generated templates, simulating a real-world photo booth environment to gauge direct consumer preferences. Lastly, we use the trained predictor to calculate the choice probabilities of the generated templates, providing an analytical measure of their expected popularity based on historical data and consumer behavior patterns. Together, these methods offer a comprehensive evaluation of the templates’ performance by considering similarity to established popular templates, direct consumer feedback, and predictive analytics.
The evaluation methods consistently indicate that the advanced model, which integrates both internal and external consumer preferences, performs the best. The designs generated by this model are predicted to be more favored by consumers, as confirmed by the distance-based metric, real consumer feedback, and the predictor model. The external dataset plays a crucial role in this process. It provides valuable training inputs for the generative models, enabling them to incorporate new product features from external designs and address the limitation of style variations in the internal data. Additionally, the rich external preferences allow the predictor model to identify popular features from external sources, guiding the generative model to create novel and attractive designs that are absent in the internal data. Two key insights emerge from the analyses: First, relying solely on internal preferences is insufficient. Incorporating external preferences is crucial as they offer additional insights into how consumers evaluate external designs. Without this information, the generative model may fail to integrate popular features from external designs that are missing in internal ones. Second, adapting the CcGAN model to acknowledge the ordinal nature of consumer preference labels is essential; without this adaptation, the generated designs perform worse.
We further conduct a personalized design generation exercise. We randomly select an internal consumer and use the predictor model to estimate his choice probabilities, termed personalized popularity labels. We use these personalized labels to re-train the general advanced model. This personalized CcGAN is used to generate new templates tailored to the consumer’s preferences. The personalized CcGAN outperforms the general CcGANs, demonstrating our model’s capability of conducting personalized design generation.
Our study bridges quantitative marketing research with generative AI advancements. While marketing literature has focused on understanding consumer preferences to guide human designers (Hauser and Clausing, 1988; Griffin and Hauser, 1993), and computer science has developed generative capabilities without systematic preference integration (Brock et al., 2018), our work merges these approaches to create a framework that is both generative and preference-aware. This integration enables product design processes to be data-driven, automated, and consumer-centric, establishing a new paradigm that transforms how firms make product design decisions.
@@
Managerially, our work offers an automated product design solution that is particularly beneficial for smaller start-up firms facing cold-start issues and budget constraints. Firms typically possess internal data regarding their existing product designs and consumer choices. Our framework allows them to systematically, scalably, and cost-effectively integrate product design ideas and consumer preferences from external websites. It also dynamically evolves as more internal and external data are incorporated, enabling firms to quickly adapt to shifts in consumer demand.
@@
The framework transforms traditional product design in three key ways: First, it provides a superior alternative to industry practices relying solely on pure generative models, which ignore consumer tastes during design generation, necessitating ex post facto screening of unpopular designs. Our method generates designs preferred by consumers ex ante, reducing time and resources spent on less appealing options. Second, it offers an effective solution for product design ideation, moving beyond reliance on human designers with limited capacity. By leveraging UGC, it captures real-time preferences from a broad user base. Third, it provides a cost-effective way to obtain consumer evaluations of new designs by using UGC as a source of self-revealed and dynamic consumer preferences, contrasting with traditional costly surveys. This approach is applicable in other settings, such as video meeting backgrounds, where consumer preference information and external resources are abundant.
Literature Review
Our paper contributes to the literature on the usage of UGC in marketing. Most of these studies were descriptive and focused on analyzing UGC to extract information and gain insights for firm operations. Some of them have analyzed textual content. For example, Netzer et al., (2012) extracted brand associations from UGC on forums. Culotta and Cutler, (2016) extracted the social connections of brands on Twitter. Other works analyzed visual content (Zhang and Luo,, 2023; Zhang et al.,, 2017; Pavlov and Mizik,, 2019) . For instance, Liu et al., (2020) analyzed consumer-generated images on social media to study the brand perceptions of consumers. We contribute to the literature by going one step further, from descriptive to prescriptive work. Our generative model not only extracts information from unstructured UGC but also uses this information to make unstructured product design decisions for firms.
We contribute to the literature on deep learning models in marketing. Most studies in this area use unstructured data to generate structured firm insights. For instance, Liu et al., (2020) collected unstructured image data from social media and used a deep convolutional neural network (CNN) to extract and predict consumer perceptions of brand attributes. Our work differs in that we employ unstructured data to directly generate unstructured firm decisions (i.e., product designs as images) . Our proposed framework goes beyond the descriptive task and provides prescriptive solutions for the product design decisions of firms.
We also contribute to the nascent literature on deep generative models in marketing. A few works have used deep generative models to create new product designs. For example, Dew et al., 2022 used multiview representation learning to design brand logos. Burnap et al., 2023 proposed a model to augment the commonly used aesthetic design process by predicting aesthetic scores and automatically generating innovative and appealing product designs; the predictor model and the generator models were separately trained. Our work differs in that we incorporate consumer preference information into the generation process so that the predictor model guides the generative model. We also introduce and adapt the CcGAN model to the marketing field, allowing us to systematically incorporate consumer preferences in an automated and large-scale manner. In addition, while previous works generated one aspect of product designs such as logos (Dew et al.,, 2022) or the aesthetic appearances of cars (Burnap et al.,, 2023), we directly design the entire product (i.e., the photo templates are the products of the firm). This setting allows us to abstract away from other design factors that might influence consumer choices.
Relationship to other generative models
Our work builds on GAN models, which offer distinct advantages over other generative models like diffusion models. While diffusion models excel in producing high-resolution images with fine details, they are more suited for generating surreal scenes and struggle with practical imagery commonly used for information sharing and creation. [^3] https://openai.com/index/introducing-4o-image-generation/ Additionally, diffusion models are computationally intensive and slower due to their iterative denoising process. State-of-the-art diffusion models require at least 10 seconds to generate one image and rely on 2 to 3 enterprise-grade GPUs. Advanced models like GPT‑4o image generation, launched in March 2025, can produce high-quality realistic images but require textual prompts and take 30-60 seconds per image, with some taking up to 2 minutes. [^4] https://help.openai.com/en/articles/8932459-creating-images-in-chatgpt In contrast, GANs generate more realistic images, are less computationally demanding, simpler to implement and tune, and enable faster generation, making them more suitable for our goal of large-scale product design generation. [^5] https://www.geeksforgeeks.org/generative-adversarial-networks-gans-vs-diffusion-models/ Our trained model can generate a new design in just 0.01 seconds using a consumer-grade laptop, which is 1,000 times faster than diffusion models. In addition, considering the difference in hardware requirements, the generation cost can be 8,000 to 12,000 times lower. [^6] Our model can generate an image in 0.01 seconds using a consumer-grade laptop. The cost of using an AWS G4 instance (equivalent to a consumer-grade laptop GPU) is \$1.461 to generate 1 million images. In contrast, diffusion models need 3 enterprise-grade GPUs to generate an image in 10 seconds. The cost for 3* A100 GPUs on AWS is \$12,291 to produce 1 million images. Alternatively, generative AI services like DALL-E 2 charge \$18,000 for 1 million images, while Stable Diffusion 3 costs \$35,000 for the same amount of images. The online appendix provides more calculation details.
Another advantage of the GAN models lies in their conceptual attractiveness. Continuous diffusion models predominantly operate within non-equilibrium thermodynamic frameworks, utilizing score matching methods that progressively transform noise distributions through iterative denoising processes governed by stochastic differential equations. This paradigm inherently embraces non-equilibrium dynamics, where the generative process follows irreversible trajectories through probability density spaces. In contrast, our CcGAN architecture extends the min-max game-theoretic framework, which fundamentally aligns with the equilibrium-seeking conceptualizations prevalent in marketing and applied microeconomics. This allows our model to converge toward Nash equilibria where generator and discriminator functions achieve optimal strategic responses to each other, mirroring the equilibrium-based analytical constructs that underpin consumer preference modeling and market clearing mechanisms in microeconomics. The theoretical congruence between CcGAN’s adversarial optimization and economic equilibrium concept is further corroborated by the self-play theorem proving systems (exemplified by DeepSeek’s architecture), which demonstrate that adversarial self-competition frameworks with primitive reward-punishment signals can systematically converge toward optimal solutions through iterative strategic adaptation. This mathematical homology between adversarial generative models and strategic equilibrium-seeking economic agents provides our framework with superior interpretability within established paradigms, while maintaining competitive generative performance compared to diffusion-based alternatives.
Our paper relates to the literature on image generation incorporating human preference, which is a rapidly growing research area in computer vision. State-of-the-art models mainly use diffusion-based models, as seen in works like Xu et al., (2023) and Wallace et al., (2024). Our model differs in the following ways. First, existing works focus on text-to-image diffusion models, in which human preference is represented by prompt alignment, or human ratings on how well the generated images align with specific text prompts (Xu et al., 2023). By nature, this type of human preference differs from consumer preferences in marketing, which focus on the mapping between images and consumers’ tastes. Second, human preference in diffusion models only captures the average person’s preference and overlooks consumer diversity. For instance, in Xu et al., (2023), preference is vertical, meaning everyone agrees on image quality rankings. In contrast, marketing highlights horizontal preference, where consumers have varying tastes in image styles. Our model addresses this by capturing diverse preferences among heterogeneous consumers. Importantly, our model is well designed for personalized design generation, which we demonstrate using a personalization exercise. Third, diffusion models usually require textual prompts as inputs, limiting their ability to capture large-scale consumer preferences. Our model directly uses unstructured images from consumer data as inputs without textual prompts. Consumer preference is captured by the mapping between consumers and their chosen images in an unstructured and flexible way. Fourth, diffusion models incorporate human preference through an iterative reward feedback process, which is time-consuming. Our model, based on GANs, uses a min-max operation in training, significantly accelerating design generation without iterative denoising. It reduces both training and inference costs for companies.
Conceptual Framework
Overview of the Conceptual Framework
We propose a semi-supervised deep generative framework for systematically and cost-effectively generating consumer-preferred product designs. Our model offers three key advantages. First, it enables firms to utilize external product designs from sources like social media and image websites, benefiting startups with limited internal designs. Second, it integrates consumer preferences from a firm’s internal data to ensure generated designs align with consumer interests. Third, it incorporates external consumer preference information, allowing firms to access and leverage the abundant consumer insights available in user-generated content (UGC). To demonstrate these three advantages, we present three versions of the generative model. The first is a baseline model that uses product designs from both internal and external sources, showcasing the benefit of leveraging external designs from UGC. The second variant enhances the baseline by incorporating consumer preference data from a firm’s internal sources, illustrating how internal preferences can guide the generation of more appealing designs. The third variant is the complete model, integrating product design and consumer preference information from both internal and external sources, thereby demonstrating all three advantages of our framework. This subsection outlines the concepts and workflows of the three models, with implementation details provided in the following subsections.
@@
Following industry standards, we use GANs (Goodfellow et al.,, 2014) to generate new product designs in the baseline model. We choose GAN models over other generative models such as fully visible belief networks (FVBNs) (Frey,, 1998) and variational autoencoders (VAEs) (Kingma and Welling,, 2013; Rezende et al.,, 2014) because GANs were designed to avoid the many disadvantages associated with other generative models. As we focus on image synthesis-based tasks, we use the DCGAN (Radford et al.,, 2015), which is considered one of the most important steps forward in terms of designing and training stable GANs.
@@
We first train a predictor model to learn the internal preferences of consumers, which we incorporate into the generative process. The predictor model utilizes the historical photos taken by internal consumers as inputs. It learns a mapping between a consumer’s appearance (i.e., her facial traits) and her preferred product designs, assuming that consumers with similar appearances will prefer similar product designs. To address the high dimensionality of the target problem, we preprocess the observed consumer faces and product designs using embedding models, obtaining low-dimensional vector representations for both consumer appearances and product designs. Using these embeddings as inputs, we train a deep prediction model to capture high-level abstractions in the prediction task. We call this predictor model with internal preferences “Prediction Model 1,” as illustrated in Figure 1. The trained predictor model can be used to predict the popularity of any given product design. Specifically, for each pair consisting of a product design and an internal consumer, the trained predictor calculates the probability that this consumer will choose the corresponding product. By averaging over the selection probabilities among all internal consumers, we can obtain a popularity measure of the particular product.
@@
The trained predictor model is further integrated into the generator model in a semi-supervised deep learning framework. We employ a conditional GAN model (Mirza and Osindero,, 2014), which extends the traditional GAN by enabling the conditional generation of outputs. Intuitively, the traditional GAN can use a set of dog images to produce a new set of dog images, but one cannot instruct the model to create images of a specific type of dog. The conditional GAN allows one to condition the network on additional information such as class labels. During training, the conditional GAN takes images with labels (bulldogs, golden retrievers, etc.) as inputs and learns the differences between the images with different labels. In this way, it gains the ability to generate new images with a specific label (a specific type of dog).
The conditional GAN is well suited for our task because it allows us to incorporate consumer preferences and generate new product designs that are more likely to be chosen by consumers. In our context, we use the popularity measures derived from the predictor model as additional inputs (i.e., class labels of product designs). The predicted popularity levels guide the training procedure of the conditional GAN so that internal consumer preferences are incorporated into the generative model. Once the generative model is trained, we can instruct the model to generate new product designs with higher popularity levels. In this way, the conditional GAN can generate new designs that are more appealing to consumers.